# Calibrating Neural Networks’ parameters through Optimal Contraction in a Prediction Problem

### v1

@@ 2.1 Neural Networks in a different domain
To assure a continuous and differentiable behavior of $U$ we must impose some requirements. In the following, we assume $F()$ is Lipschitz continuous with a Lipschitz constant of 1 or less. We notice that according to Banach’s Contraction Mapping Theorem, a solution to 5 always exists and is unique when the norm of $W$ is less than one. This fact will help us mold the Statistical Model.
@@ 2.4 Existence, Uniqueness and an Iterative Method
Assume an optimal weight matrix $W^{*}$ exists and an invertible matrix $\tilde{W}$ exists such that $(\tilde{W})^{-1}W^{*}\geq 0$ and $W^{*}(\tilde{W})^{-1}\geq 0$ (entry wise). Then, define $S$ as the Subset of pairs $(U,\mu)$ where all the matrices $W(U,\mu)$ are such that $(W^{*})^{-1}W(U,\mu)\geq 0$ and $W(U,\mu)(W^{*})^{-1}\geq 0$ (entry wise).
Then, when $\theta_{W}>\kappa(G_{X,\beta})\kappa(\tilde{W})\frac{1+\beta^{\prime}\beta+\left\|Q\right\|}{\beta^{\prime}\beta\left\|Q\right\|}$ it is possible to find a small enough $\delta$ so that the iterative method given by 13 will converge to a unique fixed point in $S$.
@@
Let us use Banach’s Contraction Mapping Theorem, again. First, we notice that $S$ is not void because an optimal weight matrix exists (as Brouwer’s Fixed Point Theorem shows) and by assumption it is part of the subset. Also, if $W^{*}$ is invertible, we can assure a matrix $\tilde{W}$ that meets the conditions exist (simply choose $\tilde{W}=W^{*}$).
@@
$\left\|\begin{pmatrix}B_{1}(U_{1},\mu_{1})-B_{1}(U_{2},\mu_{2})\\ B_{2}(U_{1},\mu_{1})-B_{2}(U_{2},\mu_{2})\end{pmatrix}\right\|\leq\frac{\kappa(\tilde{W})}{\theta_{W}}\left\|\begin{pmatrix}U_{1}-U_{2}\\ \mu_{1}-\mu_{2}\end{pmatrix}\right\|$ (17)
@@
Using Lemma 5 again, we conclude that when $\theta_{W}>\kappa(G_{X,\beta})\kappa(\tilde{W})\frac{1+\beta^{\prime}\beta+\left\|Q\right\|}{\beta^{\prime}\beta\left\|Q\right\|}$ we can find a $\delta$ so that the distance between different $(U,\mu)$ pairs diminishes in $S$.
@@
This is a looser condition than the one found in theorem 1, because convergence requires smaller parameters. This is because we have found optimal parameters are surrounded by alternatives that will iteratively converge as long as the weight matrix $W^{*}$ that arises from optimal $(U,\mu)$ pairs is invertible.
@@ 3.   Constrained RNNs and Feedforward networks
    an invertible matrix $\tilde{W}$ exists, so that a subset $S$ where all the matrices $W(U,\mu)$ arising from pairs $(U,\mu)$ are such that $(\tilde{W})^{-1}W(U,\mu)\geq 0$ and $W(U,\mu)(\tilde{W})^{-1}\geq 0$. Also, parameters $\theta_{W}$ and $\theta_{V}$ are such that $\theta_{W}>\kappa(\hat{G}_{X,\beta})\kappa(\tilde{W})\frac{1+\beta^{\prime}\beta+\left\|\hat{Q}\right\|}{\beta^{\prime}\beta\left\|\hat{Q}\right\|}$
@@
Then, simple algebra shows that whenever $w_{3,5}$ and $w_{4,5}$ have opposing signs it will be impossible to find an invertible matrix $\tilde{W}$ where $(\tilde{W})^{-1}W^{*}\geq 0$ and $W^{*}(\tilde{W})^{-1}\geq 0$ (entry wise).
@@ A.4 Lemma 4
Assume an optimal weight matrix $W^{*}$ exists. Also assume a matrix $\tilde{W}$ exists, is invertible, and $\bigl{(}\tilde{W}\bigr{)}^{-1}W^{*}\geq 0$ and $W^{*}\bigl{(}\tilde{W}\bigr{)}^{-1}\geq 0$ (entry wise).
Then, define $\tilde{S}$ as the Subset of pairs $(U,\mu)$ where all the matrices $W(U,\mu)$ are such that $\bigl{(}\tilde{W}\bigr{)}^{-1}W(U,\mu)\geq 0$ and $W(U,\mu)\bigl{(}\tilde{W}\bigr{)}^{-1}\geq 0$ (entry wise).
@@
$\left\|\begin{pmatrix}F\left(U_{1}\right)W(U_{1},\mu_{1})-F\left(U_{2}\right)W(U_{2},\mu_{2})\\ \dot{F}\left(U_{1}\right)\circ(\mu_{1}W^{\prime}(U_{1},\mu_{1}))-\dot{F}\left(U_{2}\right)\circ(\mu_{2}W^{\prime}(U_{2},\mu_{2}))\end{pmatrix}\right\|\leq\frac{\kappa(\tilde{W})}{\theta_{W}}\left\|\begin{pmatrix}U_{1}-U_{2}\\ \mu_{1}-\mu_{2}\end{pmatrix}\right\|$ (A54)
@@
$\Bigl{(}F\left(U_{1}\right)W(U_{1},\mu_{1})(\tilde{W})^{-1}-F\left(U_{2}\right)W(U_{2},\mu_{2})(\tilde{W})^{-1}\Bigr{)}\tilde{W}$
@@
$\Bigl{(}\mu_{1}((\tilde{W}^{-1}W(U_{1},\mu_{1}))^{\prime}-\mu_{2}((\tilde{W})^{-1}W(U_{2},\mu_{2}))^{\prime}\Bigr{)}(\tilde{W})^{\prime}$
That is, the condition implies $W(U_{1},\mu_{1})\bigl{(}\tilde{W}\bigr{)}^{-1}$ and $W(U_{2},\mu_{2})\bigl{(}\tilde{W}\bigr{)}^{-1}$ have the same positive (entry wise) sign.
We also notice that $\left\|W(U,\mu)\>\bigl{(}\tilde{W}\bigr{)}^{-1}\right\|\leq\frac{1}{\theta_{W}}\left\|\bigl{(}\tilde{W}\bigr{)}^{-1}\right\|$, which means that $\frac{W(U,\mu)\>\bigl{(}\tilde{W}\bigr{)}^{-1}}{\frac{1}{\theta_{W}}\left\|\bigl{(}\tilde{W}\bigr{)}^{-1}\right\|}\in[0,1]$. Therefore, the pre and post multiplication of these matrices only dampens differences:
$\begin{pmatrix}\left\|F\left(U_{1}\right)W(U_{1},\mu_{1})-F\left(U_{2}\right)W(U_{2},\mu_{2})\right\|\\ \left\|\dot{F}\left(U_{1}\right)\circ(\mu_{1}W^{\prime}(U_{1},\mu_{1}))-\dot{F}\left(U_{2}\right)\circ(\mu_{2}W^{\prime}(U_{2},\mu_{2}))\right\|\end{pmatrix}\leq\frac{\left\|\tilde{W}\right\|\left\|\bigl{(}\tilde{W}\bigr{)}^{-1}\right\|}{\theta_{W}}\begin{pmatrix}\left\|U_{1}-U_{2}\right\|\\ \left\|\mu_{1}-\mu_{2}\right\|\end{pmatrix}$ (A56)
@@
$\left\|\begin{pmatrix}F\left(U_{1}\right)W(U_{1},\mu_{1})-F\left(U_{2}\right)W(U_{2},\mu_{2})\\ \dot{F}\left(U_{1}\right)\circ(\mu_{1}W^{\prime}(U_{1},\mu_{1}))-\dot{F}\left(U_{2}\right)\circ(\mu_{2}W^{\prime}(U_{2},\mu_{2}))\end{pmatrix}\right\|\leq\frac{\kappa\bigl{(}\tilde{W}\bigr{)}}{\theta_{W}}\left\|\begin{pmatrix}U_{1}-U_{2}\\ \mu_{1}-\mu_{2}\end{pmatrix}\right\|$ (A57)

### v2

@@ 2.1 Neural Networks in a different domain
To assure a continuous and differentiable behavior of $U$ we must impose some requirements. In the following, we assume $\frac{df_{i}}{u_{i}}\in[0,1]$ so that $F()$ is Lipschitz continuous with a Lipschitz constant of 1 or less. We notice that according to Banach’s Contraction Mapping Theorem, a solution to 5 always exists and is unique when the norm of $W$ is less than one. This fact will help us mold the Statistical Model.
@@ 2.4 Existence, Uniqueness and an Iterative Method
Assume an optimal weight matrix $W^{*}$ exists and an invertible matrix $\Omega$ exists such that $(\Omega W^{*}\geq 0$ and $W^{*}\Omega\geq 0$ (entry wise). Then, define $S$ as the Subset of pairs $(U,\mu)$ where all the matrices $W(U,\mu)$ are such that $\Omega W(U,\mu)\geq 0$ and $W(U,\mu)\Omega\geq 0$ (entry wise).
Then, when $\theta_{W}>\kappa(G_{X,\beta})\kappa(\Omega)\frac{1+\beta^{\prime}\beta+\left\|Q\right\|}{\beta^{\prime}\beta\left\|Q\right\|}$ it is possible to find a small enough $\delta$ so that the iterative method given by 13 will converge to a unique fixed point in $S$.
@@
Let us use Banach’s Contraction Mapping Theorem, again. First, we notice that $S$ is not void because an optimal weight matrix exists (as Brouwer’s Fixed Point Theorem shows) and by assumption it is part of the subset. Also, if $W^{*}$ is invertible, we can assure a matrix $\Omega$ that meets the conditions exist (simply choose $\Omega=(W^{*})^{-1}$).
@@
$\left\|\begin{pmatrix}B_{1}(U_{1},\mu_{1})-B_{1}(U_{2},\mu_{2})\\ B_{2}(U_{1},\mu_{1})-B_{2}(U_{2},\mu_{2})\end{pmatrix}\right\|\leq\frac{\kappa(\Omega)}{\theta_{W}}\left\|\begin{pmatrix}U_{1}-U_{2}\\ \mu_{1}-\mu_{2}\end{pmatrix}\right\|$ (17)
@@
Using Lemma 5 again, we conclude that when $\theta_{W}>\kappa(G_{X,\beta})\kappa(\Omega)\frac{1+\beta^{\prime}\beta+\left\|Q\right\|}{\beta^{\prime}\beta\left\|Q\right\|}$ we can find a $\delta$ so that the distance between different $(U,\mu)$ pairs diminishes in $S$.
@@
This is a looser condition than the one found in theorem 1, because convergence requires smaller parameters. This is because we have found optimal parameters are surrounded by alternatives that will iteratively converge as long as there is a matrix $\Omega$ that is invertible.
@@ 3.   Constrained RNNs and Feedforward networks
    an invertible matrix $\Omega$ exists, so that a subset $S$ where all the matrices $W(U,\mu)$ arising from pairs $(U,\mu)$ are such that $\Omega W(U,\mu)\geq 0$ and $W(U,\mu)\Omega\geq 0$. Also, parameters $\theta_{W}$ and $\theta_{V}$ are such that $\theta_{W}>\kappa(\hat{G}_{X,\beta})\kappa(\Omega)\frac{1+\beta^{\prime}\beta+\left\|\hat{Q}\right\|}{\beta^{\prime}\beta\left\|\hat{Q}\right\|}$
@@
Then, simple algebra shows that whenever $w_{3,5}$ and $w_{4,5}$ have opposing signs it will be impossible to find an invertible matrix $\Omega$ where $(\Omega W^{*}\geq 0$ and $W^{*}\Omega\geq 0$ (entry wise).
@@ A.4 Lemma 4
Assume an optimal weight matrix $W^{*}$ exists. Also assume a matrix $\Omega$ exists, is invertible, and $\Omega W^{*}\geq 0$ and $W^{*}\Omega\geq 0$ (entry wise).
Then, define $\tilde{S}$ as the Subset of pairs $(U,\mu)$ where all the matrices $W(U,\mu)$ are such that $\Omega W(U,\mu)\geq 0$ and $W(U,\mu)\Omega\geq 0$ (entry wise).
@@
$\left\|\begin{pmatrix}F\left(U_{1}\right)W(U_{1},\mu_{1})-F\left(U_{2}\right)W(U_{2},\mu_{2})\\ \dot{F}\left(U_{1}\right)\circ(\mu_{1}W^{\prime}(U_{1},\mu_{1}))-\dot{F}\left(U_{2}\right)\circ(\mu_{2}W^{\prime}(U_{2},\mu_{2}))\end{pmatrix}\right\|\leq\frac{\kappa(\Omega)}{\theta_{W}}\left\|\begin{pmatrix}U_{1}-U_{2}\\ \mu_{1}-\mu_{2}\end{pmatrix}\right\|$ (A54)
@@
$\Bigl{(}F\left(U_{1}\right)W(U_{1},\mu_{1})\Omega-F\left(U_{2}\right)W(U_{2},\mu_{2})\Omega\Bigr{)}(\Omega)^{-1}$
@@
$\Bigl{(}\mu_{1}(\Omega W(U_{1},\mu_{1}))^{\prime}-\mu_{2}(\Omega W(U_{2},\mu_{2}))^{\prime}\Bigr{)}(\Omega)^{-T}$
That is, the condition implies $W(U_{1},\mu_{1})\Omega$ and $W(U_{2},\mu_{2})\Omega$ have the same positive (entry wise) sign.
We also notice that $\left\|W(U,\mu)\>\Omega\right\|\leq\frac{1}{\theta_{W}}\left\|\Omega\right\|$, which means that $\frac{W(U,\mu)\>\Omega}{\frac{1}{\theta_{W}}\left\|\Omega\right\|}\in[0,1]$. Therefore, the pre and post multiplication of these matrices only dampens differences:
$\begin{pmatrix}\left\|F\left(U_{1}\right)W(U_{1},\mu_{1})-F\left(U_{2}\right)W(U_{2},\mu_{2})\right\|\\ \left\|\dot{F}\left(U_{1}\right)\circ(\mu_{1}W^{\prime}(U_{1},\mu_{1}))-\dot{F}\left(U_{2}\right)\circ(\mu_{2}W^{\prime}(U_{2},\mu_{2}))\right\|\end{pmatrix}\leq\frac{\left\|(\Omega)^{-1}\right\|\left\|\Omega\right\|}{\theta_{W}}\begin{pmatrix}\left\|U_{1}-U_{2}\right\|\\ \left\|\mu_{1}-\mu_{2}\right\|\end{pmatrix}$ (A56)
@@
$\left\|\begin{pmatrix}F\left(U_{1}\right)W(U_{1},\mu_{1})-F\left(U_{2}\right)W(U_{2},\mu_{2})\\ \dot{F}\left(U_{1}\right)\circ(\mu_{1}W^{\prime}(U_{1},\mu_{1}))-\dot{F}\left(U_{2}\right)\circ(\mu_{2}W^{\prime}(U_{2},\mu_{2}))\end{pmatrix}\right\|\leq\frac{\kappa\bigl{(}\Omega\bigr{)}}{\theta_{W}}\left\|\begin{pmatrix}U_{1}-U_{2}\\ \mu_{1}-\mu_{2}\end{pmatrix}\right\|$ (A57)